OpenStudy (habbababba1):

If f(x) = 3x – 2 and g(x) = 6 – 4x, find f(x) + g(x). A. 7 + 4x B. 4 – x C. –x – 4 D. –4 – 7x

OpenStudy (amorfide):

if I had 3=2+1, and 4=2+2, then 3+4=(2+1) + (2+2) so in your case, f(x)=3x-2, g(x)=6-4x, what is f(x)+g(x)